# Reassessing Collaborative Writing Theories and Frameworks in the Age of LLMs: What Still Applies and What We Must Leave Behind

### v2

@@ 3. Survey Method
Drawing on the principal search system selections by Moher et al. (Moher et al., 2009), we selected 2 databases (Web of Science and the ACM Digital Library) and used the queries in Table 1. Because the selection of the ACM Digital Library was primarily to address RQ2, we added an additional filtering for LLM related papers, which was not necessary for Web of Science due to its broader, non-HCI-specific scope. Given that collaborative writing has been studied for decades, we did not limit by publication year.
@@
These queries resulted in a total of 3,067 entries (2,965 from Web of Science, 102 from ACM). After removing 7 duplicates, we applied the inclusion/exclusion criteria in Appendix A to the 3,060 entries based on the titles and abstracts. This left us with 65 entries – we lightly reviewed the full text, and screened with the predefined QA checklist in Appendix A. We applied a threshold of for the QA checklist (i.e. we only retained papers that achieved a full score of ), leaving us with 22 entries. Finally, we added 6 papers from backward referencing, resulting in a total of 28 papers for the full review. For each paper, we extracted the collaborative writing theory or framework that the work introduced or built on top of. The results of the data extraction are presented in Table LABEL:tab:papers.
@@
To ensure an objective process, two reviewers (the first and second author) took the same 15 randomly selected entries and applied the inclusion/exclusion criteria screening to check agreement levels. This sample screening not only revealed substantial agreement between the 2 reviewers (Cohen’s ), it also helped to add specificity to the exclusion criteria, which the first author applied to the rest of the entries. The first author conducted the full-text screening and the data extraction alone, with the second and third authors helping to resolve any conflicts and achieve consensus.
@@ 4.1.1. Always available for action and do not judge
As obvious as this may sound, LLMs need no rest. They are “always” available (except due to technical downtime, etc.), and there is no need for social hesitation or awkwardness when asking for help. As Gero et al. (Gero et al., 2023)’s work revealed, writers feel there is no need to work hard to establish or maintain a good relationship with AI agents, which frees them from the worry of being judged. While this may seem contrary to the CASA paradigm, Ede and Lunsford (Ede and Lunsford, 2006) interestingly point out that in the most ideal form of collaborative writing, you should “let everything show - the stupidities, the backsliding and illogicalities, and everything else that goes into writing a text - as well as to be open to those things in the other person, and to the other person’s strengths”. This suggests that, if used appropriately, the lack of social norm could paradoxically be what allows for exceptionally effective collaboration, be it for human-human or human-AI.
In other words, writers are able to ask from LLMs much more than what they would request from humans, without worrying about social norms. For example, Laban et al. (Laban et al., 2024) experimented with executable edits which allowed authors to edit faster than simple chat interactions. There is much to borrow from AI-powered code editors such as Cursor222See https://www.cursor.com or Github Copilot333See https://github.com/features/copilot in this aspect, where it is common for suggestions to be applied to the codebase with a click, beyond suggestions via chat.
@@ 4.1.2. No need for consensus building
Within the various definitions for collaborative writing we found in our review, some necessitate that all decisions be “made by consensus” (Lay and Karis, 1991). Ede and Lunsford (Ede and Lunsford, 2006) also propose the dialogic mode of writing (as opposed to the hierarchical mode), where all participants have equal input, shared authorship, and is suited for creativity, requiring consensus and compromise. Lowry (Lowry et al., 2004) labels this as “reactive writing”. Since AI tools have no opinions, values, or goals, we believe there is no need to consider building consensus with AI. However, there is value in human writers making their values and goals clear to an AI writing tool, and for human co-authors to do so with potential support from an AI writing tool, as we will discuss in our findings.
@@ 4.1.4. Distrust is encouraged
As meaningless as it may be to state, machines don’t care. Therefore, a certain level of distrust and skepticism towards AI as a co-writer is, ironically, encouraged. This is clearly different from human-human collaborative writing, where the establishment of trust and camaraderie is seen as a major component for success (Ens et al., [n. d.]; Lowry et al., 2004). We build on the approach of Cai et al. (Cai et al., 2019a) and Gu et al. (Gu et al., 2021) in creating clinical decision support tools: instead of waiting for AI to be good enough to be “fully automatable one day” (Gu et al., 2021), we should accept its “imperfection” and move towards creating tools with careful HCI considerations, which, in turn, would lead to better calibrated trust (Jacovi et al., 2021) and utility while experiencing less workload (Cai et al., 2019a).
@@ 4.2. Seven insights and design implications from legacy theories/frameworks
| Theme | Design implication |
| Writing process and strategies are iterative, non-linear and interchangeable any time | Allow for edits or additional prompts at any given moment during the process of writing |
| Supporting planning has been the primary focus in collaborative writing tools | Equip users with granular revision control |
| Design novel methods to effectively prototype complex documents | |
@@
| Provide built-in verification methods for AI-generated text | |
@@
4.2.1. Writing process and strategies are iterative, non-linear and interchangeable any time
@@
Cognitively, writing processes are highly flexible and non-linear. This is one of the key insights of the Cognitive Process Theory of Writing (Flower and Hayes, 1981), which was the first to model the cognitive process of writing and has been referred to by many collaborative writing studies since. They proposed a hierarchically-organized model that primarily consists of 3 steps: (1) planning; (2) translating; and (3) reviewing, but writers “do not march through these processes in a simple 1, 2, 3 order” (Flower and Hayes, 1981). The writing process is essentially a constant iteration of creating and revising goals and sub-goals, and therefore key subprocesses such as idea generation, revision and evaluation can and should interrupt other processes at any given moment during the process of writing. More simply put: revision is not something that happens after writers do their writing. Rather, it happens “while in the midst of generating a text” (Faigley and Witte, 1981).
@@
This non-linear and interchangeable nature of the writing process is a common theme across various theories and frameworks. Ede and Lunsford (Ede and Lunsford, 2006), when explaining their proposed model of hierarchical and dialogic modes of writing, equally stress the importance of resisting the urge to treat writing as a simple dichotomy between the two: “it is inherently mixed and paradoxical”. Even prior to Google Docs, collaborative writing tools have supported parallel-partitioned writing, allowing members to view and edit each other’s work at any given moment, which is known to increase teamwork effectiveness (Lowry and Nunamaker, 2003).
Design implication #1: Allow for edits or additional prompts at any given moment during the process of writing
Even if AI co-writers become the ones that generate most of the actual text, it is still vital for collaborative writing tools to respect and support the human’s cognitive processes. While we found studies that quite literally build on the 3 steps of the Cognitive Process Theory of Writing and incorporate its non-linearity (Chakrabarty et al., 2024; Zhang et al., 2023), we feel that a key design implication here is the importance of allowing the user to seamlessly revise and test new ideas at the same speed as when they are writing themselves. That is, the user should not be forced to wait for the LLM’s text generation to complete. Rather, the generation should always run in the background and be interruptable, so that the user can make edits or request for additional prompts at any given moment during the collaborative writing process. For instance, imagine an interface where the text generation is constantly run in the background, with no explicit “generate” button. We see this to be much more in line with the human’s non-linear and interchangeable cognitive process of writing, compared to a chat based interaction where the reviews, edits and imaginations are forced to take place in a linear manner.
@@ 4.2.2. Supporting planning has been the primary focus in collaborative writing tools
Traditionally, supporting planning has been the primary focus in CW tools (De Silva, 2007; Zhang et al., 2023; Lowry and Nunamaker, 2003). Prasetyo and Bandung discuss the planning features of their e-book writing tool, but features for reviewing hardly go beyond simple annotations and comments. This is understandable, since the translation phase is mostly the humans doing their work, and revision requires the humans to read each other’s work: they were both primarily human-centric tasks. However, particularly in cases where the LLMs take the role of writer, we question the necessity of such emphasis on planning. Indeed, in Chakrabarty et al.’s study on supporting professional writers with LLMs (Chakrabarty et al., 2024), the AI’s contributions with the lowest retention rate (i.e. not kept until the final document) were those related to high-level planning such as “Give me ideas for why a dog and its owner might become estranged”, while the highest retention (i.e. kept in the final document) were those related to translation such as “What’s a better name than Suzanne”. Perhaps this is reflecting one of the core differences between human and AI writers: AI does not have a stake in the written content, and therefore it is seemingly unfruitful to plan together. The strength of the AI co-writer lies not in the meticulous refinement of the message to convey, but simply in the faster and easier translation of ideas to words, as well as rigorous revision of formatting and errors. Interestingly, this shift is not entirely new in writing with technology: studies show that undergraduate writers planned less before writing when they used a word processor than with pen and paper (Hayes, 1996; Haas, 2013). If the shift from pen and paper to word processors reduced the (presumed) need for planning, then it is natural to assume that there will be bigger impacts with the shift from manual writing to LLMs. Now that the generation of text itself can be done effortlessly within seconds, we see planning to primarily emerge as a result of the human writer reflecting on the AI-generated drafts. It is not that the value of planning is diminishing, but that AI co-writers may encourage users to get something down on paper first, then revise and plan accordingly, nudging writing more towards an empirical “trial and error” process than a fully planned out one, similar to prototyping. Buschek’s work on comparing the current human-AI collaborative writing scene with literary collage from avant-garde literature (Buschek, 2024) frames this shift elegantly: writing with LLMs is closer to the assembling of “prefabricated text”, consisting of multiple voices and sources.
Design implication #2: Equip users with granular revision control & Design novel methods to effectively prototype complex documents
@@
Shifting the focus from planning to revision, however, also comes with a cost when it comes to LLMs. LLMs are prone to sycophancy bias (Sharma et al., 2023), essentially making it a “yes man” regardless of factual truth. If so, is it also unrealistic to expect genuine critical feedback from LLMs? Hints can be drawn from the Re3 framework (Ruan et al., 2024), which proposes the taxonomy of revision in Figure 3. By dividing revision by its qualitative dimensions such as action, intent and granularity, and further subdividing each by their characteristics, we could enable users to provide more granular instructions to LLMs on the level of desired revision, such as requesting for grammar or clarity related edits for intent on sentence or subsentence granularity, but forbidding factual content updates which LLMs are notoriously known to be unreliable with. Given that the revision has long been mistakenly seen as just the “tidying-up activity” after completing a first draft (Faigley and Witte, 1981), helping the human writers to get accustomed to these more granular terminologies around revision is then perhaps the key agenda here.
We also acknowledge that the instantaneous and effortless text-generation can potentially lead to the loss of critical thinking for the human writers, as already seen with image generation tools (Wadinambiarachchi et al., 2024). If we are to blindly generate documents or emails to send to our peers without reviewing their output, then the results would certainly be catastrophic, both for our cognitive process, and for workplace collegiality and responsibility. The problem here is not the ease of text generation, but the difficulty in reviewing and iterating fully developed text. Perhaps the question we should be asking ourselves here is: how can we allow for effective prototyping of complex documents? For example, by leveraging the LLM’s semantic capabilities it is now possible for us to draft with smaller or summarized chunks of text, then move on to a fully fleshed document, similar to evolving prototypes from low-fidelity to high-fidelity. We see this to be a highly interesting and novel challenge that we welcome researchers to build further.
@@ 4.2.3. Group awareness is a key to success
While a direct application of human-human group awareness to human-AI collaboration could be considered excessive anthropomorphism, we believe that defining the human’s expectation for AI tool’s roles and boundaries prior to the writing could help mitigate under- and over-reliance, as seen in the field of Explainable AI. For instance, Cai et al (Cai et al., 2019b) found that, in clinical contexts, clinicians wished to know the training data, the medical views and the design objectives of the AI, far before engaging in the explanations of their diagnosis recommendations. As long as it is in service of improving the user’s motivation towards the task at hand, it may be worthwhile to provide a quick onboarding to align on the user’s expectations on the AI tools’s roles and capabilities prior to commencing the writing. We see Wiethof et al. (Wiethof et al., 2021)’s work to be in line with this principle: they provided users with a written self-introduction by the AI agent prior to the writing, which led to “realistic expectations”. However, their implementation also came with the literal adoption of Nass and Moon’s Social Response Theory (Nass and Moon, 2000) such as providing social identities to the agent, which we consider unnecessary anthropomorphism. Similarly, Cress and Kimmerle (Cress and Kimmerle, 2023) suggest refining the human-human group awareness to suite the more asymmetrical human-AI relationship, and suggests the need for AIs to develop theory of mind, the cognitive ability to be aware of others’ mental states. A good balance here may be to simply provide a moment for the writer to individually reflect on what support they are seeking from the AI co-writer, both to nurture the self-awareness as well as to inform the AI tool of the user’s expectations and goals, but not introduce the AI co-writer as another team mate with its own values and writing styles.

### v3

@@ 3. Survey Method
Drawing on the principal search system selections by Moher et al. (Moher et al., 2009), we selected 2 databases (Web of Science and the ACM Digital Library) and used the queries in Table 1. It is important to note that this study aims to bridge existing collaborative writing theories/frameworks (which fall under various research fields, such as rhetoric and writing, cognitive science, and Computer-Supported Cooperative Work) with AI writing studies (which are often from HCI). For this reason, we relied on the single phrase “collaborative writing” for the Web of Science query to cast an intentionally wide net - we wanted to gather as much existing work from outside the fields of HCI and computer science, new and old. At the same time, we intentionally added a filtering for LLM related keywords for the ACM library query, because the selection of the ACM library was primarily to address RQ2. This was not necessary for Web of Science due to its broader, non-HCI-specific scope. Given that collaborative writing has been studied for decades, we also did not limit by publication year.
@@
These queries resulted in a total of 3,067 entries (2,965 from Web of Science, 102 from ACM). After removing 7 duplicates, we applied the inclusion/exclusion criteria in Appendix A to the 3,060 entries based on the titles and abstracts. In addition to fictional writing, rhetoric, and authorship, we also excluded studies in the educational context, such as those related to the pedagogy of collaborative writing or second language learners. This is because, in the context of organizational workplace settings, we see such considerations to be largely irrelevant – the primary aim for such users is to simply get the task done with less time and effort. As mentioned in Section 1.1, the scope of this paper is limited to the pragmatic support of writing tasks in organizational workplace settings, such as writing emails, clinicians drafting discharge summaries, meeting participants drafting a summary to send out to their team, or academics crafting grant proposals. While educational approaches in these settings would certainly be of benefit in theory, in most contexts, it would get in the way of the users’ objectives. This left us with 65 entries – we lightly reviewed the full text, and screened with the predefined QA checklist in Appendix A. We applied a threshold of for the QA checklist (i.e. we only retained papers that achieved a full score of ), leaving us with 22 entries. Finally, we added 6 papers from backward referencing, resulting in a total of 28 papers for the full review. For each paper, we extracted the collaborative writing theory or framework that the work introduced or built on top of. The results of the data extraction are presented in Table LABEL:tab:papers.
@@
To reduce potential bias, two reviewers (the first and second author) took the same 15 randomly selected entries and applied the inclusion/exclusion criteria screening to check agreement levels. This sample screening not only revealed substantial agreement between the 2 reviewers (Cohen’s ), it also helped to add specificity to the exclusion criteria, which the first author applied to the rest of the entries. The first author conducted the full-text screening and the data extraction alone, with the second and third authors helping to resolve any conflicts and achieve consensus.
@@ 4.1.1. Always available for action and do not judge
As obvious as this may sound, LLMs need no rest. They are software, and are “always” available (except due to technical downtime, etc.). Additionally, as Gero et al. (Gero et al., 2023)’s work revealed, writers feel there is no need to establish or maintain a good relationship with AI agents, which frees them from the worry of being judged.
As a result, writers are able to ask from LLMs much more than what they would request from humans, without worrying about social norms. For example, Laban et al. (Laban et al., 2024) experimented with executable edits which allowed authors to go beyond typical chat interactions and edit faster. There is much to learn from AI-powered code editors such as Cursor222See https://www.cursor.com or Github Copilot333See https://github.com/features/copilot in this regard, where it is already common for suggestions to be applied to the codebase with a click, beyond suggestions via chat.
@@ 4.1.2. No need for consensus building
Within the various definitions for collaborative writing we found in our review, some necessitate that all decisions be “made by consensus” (Lay and Karis, 1991). Ede and Lunsford (Ede and Lunsford, 2006) also propose the dialogic mode of writing (as opposed to the hierarchical mode), where all participants have equal input, shared authorship, and is suited for creativity, requiring consensus and compromise. Lowry (Lowry et al., 2004) labels this as “reactive writing”. Since AI tools have no opinions, values, or goals, we believe there is no need to consider building consensus with AI. However, there is still value in human writers communicating their values and goals clearly to an AI writing tool, and for a team of human co-authors to share among themselves with support from an AI writing tool, as we will discuss in our findings.
@@ 4.1.4. Distrust is encouraged
As meaningless as it may be to state, machines do not care. Therefore, a certain level of distrust and skepticism towards AI as a co-writer is encouraged. Jacovi et al. (Jacovi et al., 2021) distinguish trust with trustworthiness, stating that distrust between human and AI could in fact be desirable, in conditions where the humans are aware of the AI’s non-trustworthiness. This seems to be different from human-human collaborative writing literature, where the establishment of trust and camaraderie is seen as a major component for success (Ens et al., [n. d.]; Lowry et al., 2004). We build on the approach that Cai et al. (Cai et al., 2019a) and Gu et al. (Gu et al., 2021) took in creating AI decision support tools: instead of waiting for AI to be good enough to be “fully automatable one day” (Gu et al., 2021), we should accept its “imperfection” and move towards creating tools with careful HCI considerations, which, in turn, would lead to better calibrated trust (Jacovi et al., 2021) and utility while experiencing less workload (Cai et al., 2019a).
@@ 4.2. Seven insights and design implications from legacy theories/frameworks
| Themes from human-human collaborative writing | Design implications for human-AI collaborative writing |
| Writing processes and strategies are iterative, non-linear and interchangeable any time | Shift to a “prototyping” approach of text to streamline the discontinuous planning/waiting/reviewing process |
| Supporting planning has been the primary focus in collaborative writing tools | More rigor is required for helping human writers review and revise AI generated text |
@@
4.2.1. Writing processes and strategies are iterative, non-linear and interchangeable any time
@@
Cognitively, writing processes are highly flexible and non-linear. This is one of the key insights of the Cognitive Process Theory of Writing (Flower and Hayes, 1981), which was the first to model the cognitive process of writing and has been referred to by many collaborative writing studies since. They proposed a hierarchically-organized model that primarily consists of 3 steps: (1) planning; (2) translating; and (3) reviewing, with the monitoring process functioning as an orchestrator that decides when the writer moves from one process to another (Figure 1). A key aspect of this theory is that writers “do not march through these processes in a simple 1, 2, 3 order” (Flower and Hayes, 1981) - rather, it is a constant iteration of creating goals, writing towards that goal, reading what was written, revising goals, writing towards the new goal, and so forth. Therefore, key subprocesses such as idea generation, revision and evaluation are interruptable by other processes at any given moment during the writing process. More simply put: writers revise as they write, plan as they revise, and write as they plan. Contrary to many beliefs, revision is not something that happens after writers do their writing, but rather something that happens “while in the midst of generating a text” (Faigley and Witte, 1981).
@@
This non-linear and interchangeable nature of the writing process was a common theme across various theories and frameworks. When explaining their proposed model of hierarchical and dialogic modes of writing, Ede and Lunsford (Ede and Lunsford, 2006) equally stress the importance of resisting the urge to treat writing as a simple dichotomy between the two: “it is inherently mixed and paradoxical”. Even prior to Google Docs, collaborative writing tools have supported parallel-partitioned writing, allowing members to view and edit each other’s work at any given moment, which is known to increase teamwork effectiveness (Lowry and Nunamaker, 2003).
Design implication #1: Shift to a “prototyping” approach of text to streamline the discontinuous planning/waiting/reviewing process
Even if AI co-writers become the ones that generate most of the actual text, it is natural to assume the same cognitive processes will be at play for the human writer, and this should be respected and supported by AI writing tools. However, the non-linear, interchangeable nature of the cognitive process of writing is dependent on the fact that the writers are writing the text by themselves, or at least are fully aware of the written content. In the cases where the AI generates a majority of (if not the entire) the writing, the traditional planning/translating/reviewing process essentially shifts to a planning/waiting/reviewing process, thereby making the human writer’s cognitive process discontinuous due to the waiting that occurs in between. In such cases, we believe it is crucial that the AI writing tool acts as an active enabler of the human’s cognitive process. While we found studies that build on the three stages of the Cognitive Process Theory of Writing with AI (Chakrabarty et al., 2024; Zhang et al., 2023), we argue that such literal implementations act as a hindrance to the iterative interchangeable nature of the human cognitive process. Rather, AI tools should work to (1) shorten the waiting process as much as possible and (2) accelerate the cognitive process, to allow human writers to test new ideas at a similar cognitive speed as when they are writing themselves. This is where the analogy of prototyping comes in. In a prototyping process, the speed at which you iterate is key: you build something rough, test quickly, learn what could be improved, and move on to the next cycle. We argue that AI writing tools would be more effective if they support the “prototyping” of text. For instance, imagine an interface where the text generation is intentionally provided as bite-size, digestible chunks which the human writer can read and review with little time and effort, compared to a full size paragraph. The human writer can directly reorder those chunks, add additional text or instructions for further refinements and run the generation again to retrieve another round of digestible chunks, thereby ”prototyping” the writing until they are satisfied and hit the “develop full document button” for the final version. Buschek’s work on comparing the current human-AI collaborative writing scene with literary collage from avant-garde literature (Buschek, 2024) frames this approach eloquently: writing with LLMs is very much the assembling of “prefabricated text”.
Such prototyping approaches can potentially lead to the mitigation of the loss of critical thinking (Lee et al., 2025) in human writers, which are already seen in essay writing tasks (Kosmyna et al., 2025) or with image generation tools (Wadinambiarachchi et al., 2024). If we are to blindly generate documents or emails to send to our peers hardly reviewing their output, the results would be catastrophic, both for our cognitive process, and for workplace collegiality and responsibility. We argue that the problem here is not the ease of text generation, but the difficulty in reviewing and revising fully developed text. We see this to be a highly interesting and novel challenge that we welcome researchers to build further.
@@ 4.2.2. Supporting planning has been the primary focus in collaborative writing tools
Traditionally, supporting planning has been the primary focus in CW tools (De Silva, 2007; Zhang et al., 2023; Lowry and Nunamaker, 2003). Prasetyo and Bandung discuss the planning features of their e-book writing tool in depth, but features for reviewing hardly go beyond simple annotations and comments. This is understandable, since revision was primarily a human-centric task: it requires humans to read the work that they wrote. However, in cases where the LLMs take the role of writer, we question the necessity of such emphasis on planning as a function of the tool. Indeed, in Chakrabarty et al.’s study on supporting professional writers with LLMs (Chakrabarty et al., 2024), the AI’s contributions with the lowest retention rate (i.e. not kept until the final document) were those related to high-level planning such as “Give me ideas for why a dog and its owner might become estranged”, while the highest retention (i.e. kept in the final document) were those related to translation such as “What’s a better name than Suzanne”. Perhaps this is reflecting one of the core differences between human and AI writers: AI does not have a stake in the written content, and therefore it is seemingly unfruitful to plan together. The strength of the AI co-writer lies not in the meticulous refinement of the message to convey, but simply in the faster and easier translation of ideas to words, as well as rigorous revision of formatting and errors. Interestingly, we are not entirely new to this shift: studies show that undergraduate writers planned less before writing when they used a word processor than with pen and paper (Hayes, 1996; Haas, 2013). If the shift from pen and paper to word processors reduced the (presumed) need for planning, then it is natural to assume that there will be an even larger reduction when shifting from manual writing to LLMs. That is not to say, however, that the value of planning is diminishing. Planning stays just as important, but from the standpoint of appropriate allocation of tasks (Vaccaro et al., 2024), it makes more sense for the AI to prioritize translation and reviews.
Design implication #2: More rigor is required for helping human writers review&revise AI generated text
@@
Shifting the focus from planning to revision, however, also comes with a cost when it comes to LLMs. LLMs are prone to sycophancy bias (Sharma et al., 2023), essentially making it a “yes man” regardless of factual truth. If so, is it unrealistic to expect genuine critical feedback from LLMs? Hints can be drawn from the Re3 framework (Ruan et al., 2024), which proposes the taxonomy of revision as shown in Figure 4. By dividing revision by its qualitative dimensions such as action, intent and granularity, and further subdividing each by their characteristics, we could enable users to provide more granular control over revision instructions, such as requesting for grammar or clarity related edits for intent on sentence or subsentence granularity, but forbidding factual content updates which LLMs are notoriously known to be unreliable with. Given that the revision has long been mistakenly seen as just the “tidying-up activity” after completing a first draft (Faigley and Witte, 1981), helping the human writers to get accustomed to these more granular terminologies around revision is perhaps the key agenda here.
In a similar vein, Laban et al. (Laban et al., 2024) proposed the Warn-Verify-Audit framework, where new information introduced to the document by LLMs are constantly flagged, and are automatically searched online or sent to an external auditor for manual fact checking. We see similar interesting fact-checking approaches emerging (Liu et al., 2024; Amirizaniani et al., 2024; Dietz, 2024), all of which are human-in-the-loop approaches to maximize the strengths of manual oversight and LLM automation. Such novel approaches to support human writes review and revise will eventually need to be built into human-AI collaborative writing tools, to draw out the best of human-AI performance.
@@ 4.2.3. Group awareness is a key to success
As stated in Section 1.1, we prefer to see AI simply as a tool and not a team mate. While the use of human-like metaphors are certainly of benefit in human-AI collaboration, anthropomorphizing AI as a team mate often misleads users to false expectations, which is a problem when AI simply cannot be accountable for written content (Stokel-Walker, 2023). Therefore we consider straightforward translations of human-human group awareness to human-AI collaboration (such as considering the AI’s motivations, emotional state or preferred writing styles) to be excessive anthropomorphism. However, the field of Explainable AI has also shown that uncovering the human’s expectation for AI tool’s roles and boundaries prior to its use could help mitigate under- and over-reliance. For instance, Cai et al (Cai et al., 2019b) found that, in clinical contexts, clinicians wished to know the training data, the medical views and the design objectives of the AI, far before engaging in the explanations of their diagnosis recommendations. As long as it is in service of improving the user’s motivation towards the task at hand, it may be worthwhile to provide a quick onboarding to align on the user’s expectations on the AI tool’s roles and capabilities, prior to commencing the writing. We see Wiethof et al. (Wiethof et al., 2021)’s work to be in line with this principle: they provided users with a written self-introduction by the AI agent prior to the writing, which led to “realistic expectations”. However, their implementation also came with a rather literal adoption of Nass and Moon’s Social Response Theory (Nass and Moon, 2000) by providing social identities to the agent, which we see to be unnecessary anthropomorphism. Similarly, Cress and Kimmerle (Cress and Kimmerle, 2023) suggest refining human-human group awareness to suite the more asymmetrical human-AI relationship, and propose AIs to develop theory of mind, the cognitive ability to be aware of others’ mental states. A good balance here may be to simply provide a moment for the writer to individually reflect on what support they are seeking from the AI co-writer, both to nurture their own self-awareness as well as to inform the AI tool of the user’s expectations and goals, but stay away from introducing the AI co-writer as another team mate with its own values and writing styles.